OpenStudy (olivya999):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (olivya999):

A) y − 5 = 3(x − 1) B) y − 2 = −3(x − 0) C) y + 2 = 3(x + 1) D) y + 5 = −3(x + 1)

OpenStudy (olivya999):

@Preetha @jigglypuff314 @Jaynator495 @ganeshie8

OpenStudy (simplymarie_x):

First you need to find the slope of these two lines. Do you understand how to do that?

OpenStudy (olivya999):

yes, the slope is 3

OpenStudy (simplymarie_x):

Okay so then using (1,5). The point slope formula is y – y1 = m(x – x1) y - 5 = 3(x-1) is your answer

OpenStudy (simplymarie_x):

Sorry, my internet went out which was why I was slow :)

OpenStudy (simplymarie_x):

No prob!